# Motion Tracks: A Unified Representation for Human-Robot Transfer in Few-Shot Imitation Learning

### v1

@@ Abstract
Teaching robots to autonomously complete everyday tasks remains a challenge. Imitation Learning (IL) is a powerful approach that imbues robots with skills via demonstrations, but is limited by the labor-intensive process of collecting teleoperated robot data. Human videos offer a scalable alternative, but it remains difficult to directly train IL policies from them due to the lack of robot action labels. To address this, we propose to represent actions as short-horizon 2D trajectories on an image. These actions, or motion tracks, capture the predicted direction of motion for either human hands or robot end-effectors. We instantiate an IL policy called Motion Track Policy (MT-$\pi$) which receives image observations and outputs motion tracks as actions. By leveraging this unified, cross-embodiment action space, MT-$\pi$ completes tasks with high success given just minutes of human video and limited additional robot demonstrations. At test time, we predict motion tracks from two camera views, recovering 6DoF trajectories via multi-view synthesis. MT-$\pi$ achieves an average success rate of 86.5% across 4 real-world tasks, outperforming state-of-the-art IL baselines which do not leverage human data or our action space by 40%, and generalizes to scenarios seen only in human videos. Code and videos are available on our website (https://portal-cornell.github.io/motion_track_policy/).
Figure 1: Motion Track Policy (MT-$\pi$ Overview): Left: Using 10 minutes of human video and a small set of robot demonstrations, we train MT-$\pi$ to take a third-person image observation ($I_{t}$) and output motion tracks ($\hat{s}$), a cross-embodiment action space of 2D trajectories in image space representing manipulator movement. Right: At test-time, we predict $\hat{s}$ along with grasps ($\hat{g}$) from 2 camera views (1 shown) and recover full 6DoF robot actions for execution.
@@ I Introduction
An alternative approach is robot learning from passive observations, such as videos of humans demonstrating desired tasks. Human video data is far easier to collect at scale, with existing datasets providing thousands of hours of demonstrations  [11, 12]. However, these human videos lack robot action labels that are necessary for training imitation learning policies, presenting a challenge in transferring knowledge from human videos to robot policies.
@@
Our key insight is that, despite the embodiment gap between human hands and robot end-effectors, we can unify their action spaces by projecting their movements onto the image plane as 2D trajectories. We train an IL policy that takes image observations as input and predicts actions as motion tracks, or short-horizon 2D trajectories in image space indicating the predicted direction of motion for points on a human hand or robot end-effector. This reframing makes action prediction compatible with both human hands and robot end-effectors. Our approach requires minimal data: approximately 10 minutes of human video and a few dozen robot demonstrations. We obtain ground-truth motion tracks using state-of-the-art hand-tracking [22] for human observations and forward kinematics with known camera extrinsics for robot demonstrations. At test time, we predict motion tracks from two camera views and use multi-view geometry to reconstruct executable 6DoF end-effector trajectories in 3D.
@@ II Related Work
Imitation Learning: Imitation learning (IL) is a framework where an agent learns to mimic expert behavior by observing demonstrations [23, 1, 24]. This paradigm has been widely applied in robotics, where policies are trained to map sensory inputs to motor actions [25, 26, 27]. Recent advancements in policy architectures, such as Diffusion Policy [3] and Action Chunking Transformer (ACT) [4], have significantly improved the real-world applicability of IL. However, these approaches still often require a significant amount of teleoperated data to achieve decent performance.
@@
Sample-Efficient State-Action Representations: To address the issue of sample inefficiency in IL, recent works propose leveraging alternative input modalities beyond images, such as point clouds [34, 35, 36], radiance fields [37, 38], or voxel grids [39, 40]. These representations better capture 3D geometric features in visual scenes, which can lead to decent policies from fewer demonstrations. However, these approaches are only compatible with robot-only data consisting of RGB-D observations, along with calibrated camera extrinsics. Consequently, images remain a prevalent choice for policy inputs as they allow for fewer assumptions during data collection and thus are more suitable to be used with cross-embodiment data like human video.
@@
One effective strategy for methods that take images as input is to additionally condition on 2D representations like optical flow [41, 40, 42], object segmentation masks [43, 44], bounding boxes [32], or pixelwise object tracks [45, 17, 18]. While these methods focus on reparameterizing the inputs to IL policies, the action space remains rooted in robot proprioceptive states. This formulation limits the policy to learning actions only present in the teleoperated data. We instead propose an embodiment-agnostic, image-based action space which encourages behaviors captured in robot data as well as those that may only be present in human videos.
Learning from Human Video
II-A Pretrained Representations on Internet-Scale Data
@@
Given the challenges associated with collecting teleoperated robot data, human video is a popular alternative source of data that is easier to collect and more widely available. Several approaches use large-scale internet human video datasets  [11, 12, 46, 47] to pretrain visual representations for robot policy learning [14, 13, 48]. However, prior works have found these pretrained representations to be brittle when used out-of-the-box or finetuned for specific tasks due to the sheer diversity of the datasets they were trained on [15, 16].
II-B Implicit Priors from Human Video
@@
An alternative approach involves collecting task-specific human video data, along with a smaller amount of teleoperated robot data, and jointly learning on the hybrid dataset [20, 21, 49]. The primary challenge with this approach is overcoming the embodiment gap, as human video does not provide robot action labels. To address this, some methods attempt to extract priors from human video that can guide robot learning, such as shared human-robot latent embeddings with which to condition robot policies [20, 50] or reward signals for reinforcement learning [21, 51, 52, 53]. However, the former methods often require a goal image, which is not always available, and the latter demand significant on-policy robot interaction.
II-C Explicit Tracking of Human Hands
@@
Recent advancements in human-hand tracking [22, 54, 55, 56] demonstrate that we can now more reliably track the explicit motion of human hands in video, instead of only extracting implicit priors. Building off of this, several recent works demonstrate few-shot visual imitation from human video alone, without requiring any robot data [57, 58, 59, 60, 61]. These methods largely assume that detected human hand poses can be directly retargeted to robot end-effectors. However, the significant morphological differences (i.e. differences in size and shape between human and robot hands) can mean that even just replaying a tracked human hand trajectory on a robot end-effector may fail to produce the desired behavior in many cases. Our work builds on these approaches by introducing a unified action space compatible with both human and robot embodiments, where actions are represented as 2D trajectories in the image plane. As the actions themselves are now aligned in a shared space, we disentangle motion differences from visual differences, allowing us to better bridge the embodiment gap and train on a hybrid of human videos and teleoperated robot data.
@@ V-A Experimental Setup
Method Human and Robot Data Wrist Camera Input 6DoF EE Delta Action Space Diffusion Backbone
MT-$\pi$
@@ V-B Key Results and Findings
Fold Cloth Fork on Plate Serve Egg Put Away Socks
@@
Figure 4: Effect of Data Distribution on Policy Success: We evaluate MT-$\pi$ against DP and ACT on the Serve Egg task (see Fig. 1), measuring policy success (green = high, red = low) subject to varying amounts of robot and human training data. While all policies improve with more robot data, collecting teleoperated demonstrations takes nearly 5x longer than human videos (70 human demos $\sim$ 10 mins., 100 robot demos $\sim$ 65 mins.). MT-$\pi$ achieves strong performance even in the low robot data regime by leveraging just $5-10$ minutes of human video, suggesting further gains are possible by fixing the amount of robot data and scaling human data alone.
@@
Close Direction DP [3] ACT [4] MT-$\pi$ (Robot Only) MT-$\pi$ (H+R)
Left (In $D_{\text{robot}}\cup D_{\text{human}}$) 20/20 17/20 20/20 20/20
Right (Only in $D_{\text{human}}$) 0/10 0/10 0/10 18/20
@@
VI Limitations and Failure Modes
While MT-$\pi$ demonstrates sample-efficiency, generalization, and reliable performance for the most part, our policy is not without failures. Namely, our policy currently makes predictions on one image at a time, and handles different viewpoints independently. This does not explicitly enforce consistency between tracks across separate views, which can lead to triangulation errors that produce imprecise actions. We try to ensure that teleoperated demonstrations are as unimodal as possible to encourage consistency in motion recovery. In the future, we can consider more explicitly enforcing viewpoint consistency via auxiliary projection/deprojection losses. Further, as a motion-centric method, our approach remains sensitive to noise in human video inputs. This sensitivity currently limits our ability to handle truly in-the-wild videos that feature drastic viewpoint shifts, egocentric motion, or quick temporal changes. While hand tracking is a fairly reliable part of our pipeline, detection of human grasps remains an open challenge. We employ a heuristic approach at present, leveraging foundation models to infer when hands and objects are in contact. Due to some imprecision in ground truth human grasp labels, our policy occasionally prematurely or imprecisely grasps objects. Nevertheless, our framework is designed with modularity in mind, allowing us to incorporate future advancements in hand perception.

### v2

@@ Abstract
Teaching robots to autonomously complete everyday tasks remains a challenge. Imitation Learning (IL) is a powerful approach that imbues robots with skills via demonstrations, but is limited by the labor-intensive process of collecting teleoperated robot data. Human videos offer a scalable alternative, but it remains difficult to directly train IL policies from them due to the lack of robot action labels. To address this, we propose to represent actions as short-horizon 2D trajectories on an image. These actions, or motion tracks, capture the predicted direction of motion for both human hands and robot end-effectors. We instantiate an IL policy called Motion Track Policy (MT-$\pi$) which receives image observations and outputs motion tracks as actions. By leveraging this unified, cross-embodiment action space, MT-$\pi$ completes tasks with high success given just minutes of human video and limited additional robot demonstrations. At test time, we predict motion tracks from two camera views, recovering 6DoF trajectories via multi-view synthesis. MT-$\pi$ achieves an average success rate of 86.5% across 4 real-world tasks, outperforming state-of-the-art IL baselines which do not leverage human data or our action space by 40%, and generalizes to scenarios seen only in human videos. Code and videos are available on our website (https://portal-cornell.github.io/motion_track_policy/).
Figure 1: Motion Track Policy (MT-$\pi$ Overview): Left: Using 10 minutes of human video and a small set of robot demonstrations, we train MT-$\pi$ to take a third-person image observation ($I_{t}$) and output motion tracks ($\hat{s}$), a cross-embodiment action space of 2D trajectories in image space representing manipulator movement. Right: At test-time, we predict $\hat{s}$ along with grasps ($\hat{g}$) from 2 camera views (1 shown) and recover full 6DoF robot actions using stereo triangulation for execution.
@@ I Introduction
An alternative approach is robot learning from passive observations, such as videos of humans demonstrating desired tasks. Human video data is far easier to collect at scale, with existing datasets providing thousands of hours of demonstrations  [11, 12]. However, these human videos lack the robot action labels that are necessary for training imitation learning policies, presenting a challenge in transferring knowledge from human videos to robot policies.
@@
Our key insight is that, despite the embodiment gap between human hands and robot end-effectors, we can unify their action spaces by projecting their movements onto the image plane as 2D trajectories. We train an IL policy that takes image observations as input and predicts actions as motion tracks, or short-horizon 2D trajectories in image space indicating the predicted direction of motion for points on a human hand or robot end-effector. This reframing makes action prediction compatible with both human hands and robot end-effectors. Our approach requires minimal data: approximately 10 minutes of human video and a few dozen robot demonstrations. We obtain ground-truth motion tracks using state-of-the-art hand-tracking [22] for human observations and forward kinematics with known camera extrinsics for robot demonstrations. At test time, we predict motion tracks from two camera views and use multi-view geometry to reconstruct 6DoF end-effector trajectories in 3D.
@@ II Related Work
II-A Imitation Learning
Imitation learning (IL) is a framework where an agent learns to mimic expert behavior by observing demonstrations [23, 1, 24]. This paradigm has been widely applied in robotics, where policies are trained to map sensory inputs to motor actions [25, 26, 27]. Recent advancements in policy architectures, such as Diffusion Policy [3] and Action Chunking Transformer (ACT) [4], have significantly improved the real-world applicability of IL. However, these approaches still often require a significant amount of teleoperated data to achieve decent performance.
@@ II-B Sample-Efficient State-Action Representations
To address the issue of sample inefficiency in IL, recent works propose leveraging alternative input modalities beyond images, such as point clouds [34, 35, 36], radiance fields [37, 38], or voxel grids [39, 40]. These representations better capture 3D geometric features in visual scenes, which can lead to decent policies from fewer demonstrations. However, these approaches are only compatible with robot-only data consisting of RGB-D observations, along with calibrated camera extrinsics. Consequently, images remain a prevalent choice for policy inputs as they allow for fewer assumptions during data collection and thus are more suitable to be used with cross-embodiment data like human video.
@@
One effective strategy for methods that take images as input is to additionally condition on 2D representations like optical flow [41, 40, 42], object segmentation masks [43, 44], bounding boxes [32], or pixelwise object tracks [45, 17, 18]. While these methods focus on reparameterizing the inputs to IL policies, the action space remains rooted in robot proprioceptive states. This formulation limits the policy to only learning actions present in the teleoperated data. We instead propose an embodiment-agnostic, image-based action space which encourages behaviors captured in robot data as well as those that may only be present in human videos.
II-C Learning from Human Video
@@
Pretrained Representations on Internet-Scale Data. Given the challenges associated with collecting teleoperated robot data, human video is a popular alternative source of data that is easier to collect and more widely available. Several approaches use large-scale human video datasets  [11, 12, 46, 47] to pretrain visual representations for robot policy learning [14, 13, 48]. However, prior works have found these pretrained representations to be brittle when used out-of-the-box or finetuned for specific tasks due to the sheer diversity of the datasets they were trained on [15, 16].
@@
Implicit Priors from Human Video An alternative approach involves collecting task-specific human video data, along with a smaller amount of teleoperated robot data, and jointly learning on the hybrid dataset [20, 21, 49]. The primary challenge with this approach is overcoming the embodiment gap, as human video does not provide robot action labels. To address this, some methods attempt to extract priors from human video that can guide robot learning, such as shared human-robot latent embeddings with which to condition robot policies [20, 50] or reward signals for reinforcement learning (RL) [21, 51, 52, 53, 54]. However, the former methods often require a goal image, which is not always available, and the latter demand significant on-policy robot interaction to train the RL agent.
@@
Explicit Tracking of Human Hands. Recent advancements in human-hand tracking [22, 55, 56, 57] demonstrate that we can now more reliably track the explicit motion of human hands in video, instead of only extracting implicit priors. Building off of this, several recent works demonstrate few-shot visual imitation from human video alone, without requiring any robot data [58, 59, 60, 61, 62, 63]. These methods largely assume that detected human hand poses can be directly retargeted to robot end-effectors. However, the significant morphological differences (i.e. differences in size and shape between human and robot hands) can mean that even just replaying a tracked human hand trajectory on a robot end-effector may fail to produce the desired behavior in many cases. Our work builds on these approaches by introducing a unified action space compatible with both human and robot embodiments, where actions are represented as 2D trajectories in the image plane. As the actions themselves are now aligned in a shared space, we disentangle motion differences from visual differences, allowing us to better bridge the embodiment gap and train on a hybrid of human videos and teleoperated robot data.
@@ V-A Experimental Setup
Method
Human and
Robot Data
Wrist
Camera Input
6DoF $E$ Delta
Action Space
Diffusion
Backbone
MT-$\pi$ × ×
@@ V-B Key Results and Findings
Fold
Cloth
Fork on
Plate
Serve
Egg
Put Away
Socks
@@
Figure 4: Effect of Data Distribution on Policy Success: We evaluate MT-$\pi$ against DP and ACT on the Serve Egg task (see Fig. 1), measuring policy success (green = high, red = low) subject to varying amounts of robot and human training data. While all policies improve with more robot data, collecting teleoperated demonstrations takes nearly 5x longer than human videos (70 human demos $\sim$ 10 mins., 100 robot demos $\sim$ 65 mins.). MT-$\pi$ achieves strong performance even in the low robot data regime by leveraging just 5105-10 minutes of human video, suggesting further gains are possible by fixing the amount of robot data and scaling human data alone.
@@
VI Limitations and Failure Modes
While MT-$\pi$ demonstrates sample-efficiency, generalization, and reliable performance for the most part, our policy is not without failures. Namely, our policy currently makes predictions on one image at a time, and handles different viewpoints independently. This does not explicitly enforce consistency between tracks across separate views, which can lead to triangulation errors that produce imprecise actions. We try to ensure that teleoperated demonstrations are as unimodal as possible to encourage consistency in motion recovery. In the future, we can consider more explicitly enforcing viewpoint consistency via auxiliary projection/deprojection losses. Further, as a motion-centric method, our approach remains sensitive to noise in human video inputs. This sensitivity currently limits our ability to handle truly in-the-wild videos that feature drastic viewpoint shifts, egocentric motion, quick temporal changes, or occlusion of hands. While hand tracking is a fairly reliable part of our pipeline, detection of human grasps remains an open challenge. We employ a heuristic approach at present, leveraging foundation models to infer when hands and objects are in contact. Due to some imprecision in ground truth human grasp labels, our policy occasionally prematurely or imprecisely grasps objects. Nevertheless, our framework is designed with modularity in mind, allowing us to incorporate future advancements in hand perception.
Close Direction
DP [3]
ACT [4]
MT-$\pi$
(Robot Only)
MT-$\pi$ (H+R)
Left
(In $D_{\text{robot}}\cup D_{\text{human}}$)
20/20 17/20 20/20 20/20
Right
(Only in $D_{\text{human}}$)
0/10 0/10 0/10 18/20
@@ VII Discussion
VIII Acknowledgements
This work was supported in part by funds from the NSF Awards #2327973, #2006388, and #2312956, the Office of Naval Research under ONR #N00014-22-1-2293, the Google Faculty Research Award, as well as the OpenAI Superalignment Grant. Priya Sundaresan is supported by an NSF GRFP. We would like to thank Zi-ang Cao for their helpful feedback and suggestions.